ganeshie8 (ganeshie8):

show that \[\large \sum\limits_{b=1}^{\infty}\sum\limits_{a=1}^{\infty}\dfrac{1}{a^2-b^2} = \dfrac{\pi^2}{8}\] \(a\ne b\)

OpenStudy (kainui):

=D This looks fun!

ganeshie8 (ganeshie8):

Indeed it is xD my first insticts are to try partial fractions but i have no clue how to proceed really

OpenStudy (kainui):

Difference of squares seems like my first instinct, cause that'll look kind of like a geometric series hmmm

ganeshie8 (ganeshie8):

ive updated the qestion slightly..

ganeshie8 (ganeshie8):

\[\large \sum\limits_{b=1}^{\infty}\sum\limits_{a=1}^{\infty}\dfrac{1}{a^2-b^2} =\sum\limits_{b=1}^{\infty} \left(\dfrac{1}{2b}\sum\limits_{a=1}^{\infty}\dfrac{1}{a-b}-\dfrac{1}{a+b}\right)\]

ganeshie8 (ganeshie8):

idk if that telescopes ever but thats as far as i can get on my own xD

OpenStudy (anonymous):

Summation is a linear operator so you need to use partial fractions

OpenStudy (anonymous):

Sorry thought your reply is what you had to show, let me rethink lol

OpenStudy (anonymous):

The \(a\) sum apparently has a closed form, \(\dfrac{\tan{b\pi}-b\pi}{b\tan{b\pi}}\).

OpenStudy (kainui):

Wait the question has to be wrong since a=1 and b=1 is a term.

ganeshie8 (ganeshie8):

it is given as a caveat that the sum doesn't include \(a = b\) terms

OpenStudy (kainui):

Oh sorry I see, in the question the latex sometimes breaks for the first question so I sort of missed that.

ganeshie8 (ganeshie8):

from the s&g formula, \(\tan(b\pi) = 0\) for all \(b \ge 1\)... denominator becomes 0 hmm

OpenStudy (kainui):

Hmm maybe I'm wrong, but it seems like the sum should be zero, since there will always be one matching term, like a=2 and b=1 has an a=1 and b=2 term to cancel it out. \[\Large \frac{1}{2^2-1^2}+\frac{1}{1^2-2^2}=\frac{1}{3}-\frac{1}{3} = 0\]

OpenStudy (anonymous):

Maybe there's a way to rewrite the summation as a double integral?

OpenStudy (dan815):

how about looking at the continous version of it

OpenStudy (dan815):